okay i've been given a question y=(x+4)^3 and it is written in f(g(x)) form. It asks to fins f(x) and g(x) how do I find that?
Answer by stanbon(75887)

okay i've been given a question y=(x+4)^3 and it is written in f(g(x)) form. It asks to fins f(x) and g(x) how do I find that?
-------------------------------------------------
If f(g(x)) = (x+4)^3
---
g(x) = x+4
f(x) = x^3
---------------
f(g(x)) = f(x+4) = (x+4)^3
===============================
Cheers,
stan H.